It is generall a good idea to gain an understanding of the "size" of units. consider the following objects and calculate their kinetic energy.

A ladybug weighing 35.1 milligrams flies by your head at 3.75 km/h. Number A 6.35-kg bowling ball slides (not rolls) down an alley at 22.7 km/h. Number A car weighing 1080-kg is moving at a speed of 39.5 km/h. Number Based on your results above, which of the following scenarios below likely describe an object possessing one joule of kinetic energy?
Answer #1

1 km/h =1000/3600 m/s =0.27777 m/s

1.

Kinetic energy is given by

K=(1/2)MV2=(1/2)*(35.1*10-3)(3.75*0.27777)2

K=0.019 J

2.

K=(1/2)*6.35*(22.7*(0.27777))2

K=126.2 J

3.

K=(1/2)*1080*(39.5*0.27777)2

K=65010 J